# Synergizing Foundation Models and Federated Learning: A Survey

### v1

@@ 4.1.1 Parameter-Efficient Fine-Tuning
Federated Parameter-Efficient Fine-Tuning (FedPEFT), originating from the fine-tuning practices of FMs Lester et al. (2021); Hu et al. (2022); Li and Liang (2021), is a suite of techniques designed to reduce both the computational load and the associated communication overheads Malaviya et al. (2023); Woisetschläger et al. (2024). In alignment with existing FM fine-tuning taxonomies Lialin et al. (2023); Ding et al. (2023), we present FedPEFT methods in three categories: selective methods, additive methods, and reparameterization-based methods.
Selective Methods
@@
Selective methods fine-tune a small subset of the parameters, leaving the majority unchanged. In the field of LLMs, a prominent example of such methods is BitFit Ben Zaken et al. (2022), which only fine-tunes the bias terms. BitFit has inspired a series of studies in FedPEFT Bu et al. (2022); Sun et al. (2022a); Zhang et al. (2023f), demonstrating the superior communication efficiency of only updating the bias terms while still achieving competitive performance. More sophisticated methods strive to find sparse subnetworks for partial fine-tuning. Among them, various methods Seo et al. (2021); Li et al. (2021a); Tamirisa et al. (2024) advocate for the Lottery Ticket Hypothesis (LTH) Frankle and Carbin (2019), positing that a dense network contains many subnetworks whose inference capabilities are as accurate as that of the original network. FedSelect Tamirisa et al. (2024) is a representative method that encourages clients to find optimal subnetworks based on LTH and continually fine-tunes these derived subnetworks to encapsulate local knowledge. As another important aspect, RaFFM Yu et al. (2023c) proposes to prioritize specialized salient parameters by ranking them using salience evaluation metrics such as the $\ell_{1}$ and $\ell_{2}$ norms.
Additive Methods
@@
Instead of fine-tuning a subset of model parameters, additive methods incorporate lightweight trainable blocks into frozen FMs and tune the additional parameters for model adaptation. These methods not only enhance computational and communicational efficiency but also introduce an extra benefit: personalization Lu et al. (2023a), i.e., the integration of these supplementary parameters allows for the customization of heterogeneous models tailored to specific local data characteristics or user preferences. Key branches within additive methods include adapter tuning and prompt tuning. Adapter tuning integrates small-scale neural networks (known as “adapters”) into the pre-trained models Houlsby et al. (2019); Hu et al. (2022). On the other hand, prompt tuning incorporates trainable task-specific continuous prompt vectors at the input layer Liu et al. (2023a); Dong et al. (2023). More details on these methods are provided in Appendix A.
Reparameterization-based Methods
@@
The hypothesis behind reparameterization-based methods is that fine-tuning adaptations can be re-parameterized into optimization within low-rank subspaces Aghajanyan et al. (2021). Low-Rank Adaptation (LoRA) Hu et al. (2022), as a popular PEFT method from the area of LLMs, reduces the number of trainable parameters for downstream tasks by representing the weight updates with two smaller matrices (called update matrices) through low-rank decomposition Ding et al. (2023). When optimizing a parameter matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, the update equation can be written as: $\mathbf{W}\leftarrow\mathbf{W}+\Delta\mathbf{W}$. The core idea of LoRA is to freeze the original matrix $\mathbf{W}$ while approximating the parameter update $\Delta\mathbf{W}$ by low-rank decomposition matrices, i.e., $\Delta\mathbf{W}=\mathbf{A}\cdot\mathbf{B}^{\top}$, where $\mathbf{A}\in\mathbb{R}^{m\times k}$ and $\mathbf{B}\in\mathbb{R}^{n\times k}$ are the trainable parameters for task adaptation and $k\ll\min(m,n)$ is the reduced rank. The trainable parameter size is then reduced from $mn$ to $k(m+n)$. The major benefit of LoRA is that it can largely save memory and storage usage. A straightforward way to perform federated finetuning with LoRA is to train the LoRA modules $\mathbf{A}$ and $\mathbf{B}$ with homogeneous rank $k$ across all clients with standard FL such as FedAvg McMahan et al. (2017). Serval studies have shown that this method can achieve an outstanding level of trade-off between performance and communication overhead for a wide range of FMs, including language models Zhang et al. (2024b, 2023f), vision-language models Nguyen et al. (2024), and speech-to-text models Du et al. (2024).
@@
Figure 2: Taxonomy of Federated Parameter-Efficient Fine-Tuning (FedPEFT). Apart from efficiency, some methods also account for other considerations, such as data and resource heterogeneity challenges that are identified in Section 3.2 and black-box tuning (see Section 4.3).
Comparison of FedPEFT methods
@@
Figure 2 depicts the taxonomy of FedPEFT with representative methods. Note that some methods may belong to multiple overlapping categories. To compare the communication efficiency of different FedPEFT methods, Table 1 gives a brief overview of experimental evaluations from representative studies. Compared to full-model fine-tuning, FedPEFT methods only require 0.1%-30% communication overhead. We note that the differences can be attributed to several factors, including model complexity and implementation details.
4.1.2 Model Compression
Model compression refers to the techniques used to reduce the size of models, thereby improving resource efficiency Shah and Lau (2023).
@@ 4.1.2 Model Compression
Table 1: Comparison of Federated Parameter-Efficient Fine-Tuning (FedPEFT) Methods.
@@
Category Representative Work Modality Model # Full Params. # Train. Params. Training Accel. Comm. Cost Selective RaFFM Yu et al. (2023c) Txt. BERT-Large (2019) 336M 100M 6.13×6.13$\times$6.13 × 29.8% FedBF Zhang et al. (2023f) Txt. Roberta-Base (2019) 125M 0.66M $1.6\%$ Additive Adapter FedAP Zhang et al. (2023f) Txt. Roberta-Base (2019) 125M 2M $1.6\%$ FedCLIP Lu et al. (2023a) Vis.-Txt. ViT-B/32 (2020a) 150M 0.53M 3.5% FedDAT Chen et al. (2024) Vis.-Txt. ALBEF (2021b) 290M 2.86M 9.9% C2A Kim et al. (2023) Txt. DistilBERT (2020) 66M 0.06M 0.1% Fed-MNMT Liu et al. (2023d) Txt. mBART-50 (2020) 611M 8M 1.3% AdaFL Cai et al. (2023) Txt. BERT (2019) 110M 0.61M 1.63×1.63$\times$1.63 × 0.6% Prompt PromptFL Guo et al. (2023) Vis.-Txt. ViT-B/16 (2021) 87M 0.87M 2.38×2.38$\times$2.38 × 0.9% MFPT Zhao et al. (2024b) Txt. XLM-RoBERTa (2020) 270M 1.2M 0.4% FedAPT Su et al. (2024) Vis.-Txt. ViT-B/32 (2020a) 88M 2.8M 3.2% FedSP Dong et al. (2023) Txt. GPT2-XL (2019) 1.6B  [^1] M 0.5% Reparameterization-based Methods SLoRA Babakniya et al. (2023a) Txt. DistilBERT (2020) 67M 0.7M 13.47×13.47$\times$13.47 × 5.8% LP-FL Jiang et al. (2023a) Txt. BERT-Large (2019) 336M 100M 30% FedMS Wu et al. (2023c) Vis.-Txt. ViT-B/16 (2021) 87M 8.6M 10% pFedS2T Du et al. (2024) Aud. Whisper (2023) 254M 10.1M 4% FFA-LoRA Sun et al. (2024b) Txt. RoBERTa-Large (2019) 355M 0.39M 0.1%
Sparsification
Model sparsification methods reduce communication burden by only transmitting a subset of FM parameters across the network Jiang et al. (2023c). Typical methods focus on identifying and cultivating high-potential subnetworks Frankle and Carbin (2019); Tsouvalas et al. (2023).
Quantization
@@
Quantization is well-established in both the FM and FL domains Xu et al. (2024b); Reisizadeh et al. (2020), which involves decreasing the precision of floating-point parameters for mitigating the storage, computational, and communication demands. Quantization is orthogonal to other resource-efficient techniques, making it feasible to combine them for greater efficiency and flexibility Lit et al. (2022).
4.1.3 Zeroth-Order Optimization

### v2

@@ 4.1. Parameter-Efficient Fine-Tuning
Federated Parameter-Efficient Fine-Tuning (FedPEFT), originating from the fine-tuning practices of FMs (Lester et al., 2021; Hu et al., 2022; Li and Liang, 2021), is a suite of techniques designed to reduce both the computational load and the associated communication overheads (Malaviya et al., 2023; Woisetschläger et al., 2024). In alignment with existing FM fine-tuning taxonomies (Lialin et al., 2023; Ding et al., 2023), we present FedPEFT methods in three categories: selective methods, additive methods, and reparameterization-based methods.
4.1.1. Selective Methods
@@
Selective methods fine-tune a small subset of the parameters, leaving the majority unchanged. In the field of LLMs, a prominent example of such methods is BitFit (Ben Zaken et al., 2022), which only fine-tunes the bias terms. BitFit has inspired a series of studies in FedPEFT (Bu et al., 2022; Sun et al., 2022a; Zhang et al., 2023e), demonstrating the superior communication efficiency of only updating the bias terms while still achieving competitive performance. More sophisticated methods strive to find sparse subnetworks for partial fine-tuning. Among them, various methods (Seo et al., 2021; Li et al., 2021a; Tamirisa et al., 2024) advocate for the Lottery Ticket Hypothesis (LTH) (Frankle and Carbin, 2019), positing that a dense network contains many subnetworks whose inference capabilities are as accurate as that of the original network. FedSelect (Tamirisa et al., 2024) is a representative method that encourages clients to find optimal subnetworks based on LTH and continually fine-tunes these derived subnetworks to encapsulate local knowledge. As another important aspect, RaFFM (Yu et al., 2023c) proposes to prioritize specialized salient parameters by ranking them using salience evaluation metrics such as the $\ell_{1}$ and $\ell_{2}$ norms.
4.1.2. Additive Methods
@@
Instead of fine-tuning a subset of model parameters, additive methods incorporate lightweight trainable blocks into frozen FMs and tune the additional parameters for model adaptation. These methods not only enhance computational and communicational efficiency but also introduce an extra benefit: personalization (Lu et al., 2023a), i.e., the integration of these supplementary parameters allows for the customization of heterogeneous models tailored to specific local data characteristics or user preferences. Key branches within additive methods include adapter tuning and prompt tuning. Adapter tuning integrates small-scale neural networks (known as “adapters”) into the pre-trained models (Houlsby et al., 2019; Hu et al., 2022). On the other hand, prompt tuning incorporates trainable task-specific continuous prompt vectors at the input layer (Liu et al., 2023a; Dong et al., 2023).
Adapter Tuning
Adapter tuning integrates small-scale neural networks (known as “adapters”) into the pre-trained models (Houlsby et al., 2019; Hu et al., 2022). A straightforward implementation of adapter tuning is to collaboratively train a shared adapter among all clients in the FedAvg manner, as highlighted by Sun et al. (2022a). Based on FedAvg, FedCLIP (Lu et al., 2023a) incorporates an attention-based adapter for the image encoder in CLIP models (Radford et al., 2021). In the domain of multilingual machine translation, where different language pairs exhibit substantial discrepancies in data distributions, Fed-MNMT (Liu et al., 2023c) explores clustering strategies that group adapter parameters and makes inner-cluster parameters aggregation for alleviating the undesirable effect of data discrepancy. Another representative approach named C2A (Kim et al., 2023) employs hypernetworks (Ha et al., 2017) to generate client-specific adapters by conditioning on the client’s information, maximizing the utility of shared model parameters while minimizing the divergence caused by data heterogeneity.
Prompt Tuning
Prompt tuning incorporates trainable task-specific continuous prompt vectors at the input layer (Liu et al., 2023a; Dong et al., 2023). Compared to full fine-tuning, it achieves comparable performance but with $1000\times$ less parameter storage and communication (Jia et al., 2022). A variation of prompt tuning, FedPerfix (Sun et al., 2023) uses a local adapter to generate the prefixes and aggregate the original self-attention layers. Depending on target modalities, prompt tuning in current literature can be further classified into three categories:
  • Textual Prompt Tuning. Task-specific prompt embeddings are combined with the input text embeddings, which are subsequently fed into language models. These soft prompts serve as instructive contexts to influence the generation process of LLMs by steering the probability distribution of the next token (Dong et al., 2023).
  • Visual Prompt Tuning. Taking inspiration from advances in efficiently tuning LLMs, prompts are also introduced in the input space of vision models (Jia et al., 2022). Naive implementations introduce prompts at the pixel level, acting as a form of data augmentation (Li et al., 2024a). Alternatively, one could also insert the prompts as latent vectors for the first Transformer layer (Deng et al., 2024; Yang et al., 2023a). Nevertheless, an empirical study (Jia et al., 2022) has suggested that it is easier for visual prompts to learn condensed task-dependent signals in the latent input space of Transformers.
  • Textual-Visual Prompt Tuning. Unlike single-modal FMs, vision-language FMs can process and interpret both visual data and textual information, endowing them with powerful representation ability and transferability (Radford et al., 2021). Based on vision-language FMs like CLIP, textual-visual prompt tuning shows promising capabilities in FL (Guo et al., 2023), especially in cross-domain scenarios, where the model needs to generalize across varied domains and unseen classes (Qiu et al., 2024).
4.1.3. Reparameterization-based Methods
@@
The hypothesis behind reparameterization-based methods is that fine-tuning adaptations can be re-parameterized into optimization within low-rank subspaces (Aghajanyan et al., 2021). Low-Rank Adaptation (LoRA) (Hu et al., 2022), as a popular PEFT method from the area of LLMs, reduces the number of trainable parameters for downstream tasks by representing the weight updates with two smaller matrices (called update matrices) through low-rank decomposition (Ding et al., 2023). When optimizing a parameter matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, the update equation can be written as: $\mathbf{W}\leftarrow\mathbf{W}+\Delta\mathbf{W}$. The core idea of LoRA is to freeze the original matrix $\mathbf{W}$ while approximating the parameter update $\Delta\mathbf{W}$ by low-rank decomposition matrices, i.e., $\Delta\mathbf{W}=\mathbf{A}\cdot\mathbf{B}^{\top}$, where $\mathbf{A}\in\mathbb{R}^{m\times k}$ and $\mathbf{B}\in\mathbb{R}^{n\times k}$ are the trainable parameters for task adaptation and $k\ll\min(m,n)$ is the reduced rank. The trainable parameter size is then reduced from $mn$ to $k(m+n)$. The major benefit of LoRA is that it can largely save memory and storage usage. A straightforward way to perform federated fine-tuning with LoRA is to train the LoRA modules $\mathbf{A}$ and $\mathbf{B}$ with homogeneous rank $k$ across all clients with standard FL such as FedAvg (McMahan et al., 2017). Serval studies have shown that this method can achieve an outstanding level of trade-off between performance and communication overhead for a wide range of FMs, including language models (Zhang et al., 2024b, 2023e), vision-language models (Nguyen et al., 2024), and speech-to-text models (Du et al., 2024).
@@
Figure 2. Taxonomy of Federated Parameter-Efficient Fine-Tuning (FedPEFT). Apart from efficiency, some methods also account for other considerations, such as data and resource heterogeneity challenges that are identified in Section 3.2 and black-box tuning (see Section 6).
4.1.4. Comparison of FedPEFT Methods
@@
Figure 2 depicts the FedPEFT taxonomy and its representative methods. Note that some methods may belong to multiple overlapping categories. To compare the communication efficiency of different FedPEFT methods, Table 2 gives a brief overview of experimental evaluations from representative studies. Compared to full-parameter fine-tuning, FedPEFT methods only require 0.1%-30% communication overhead. We note that the differences can be attributed to several factors, including model complexity and implementation details.
4.2. Model Compression
Although FedPEFT is effective in reducing the number of trainable parameters, it does not inherently reduce the structural complexity of FMs. To further optimize the system efficiency, model compression serves as a core strategy that physically reduces the model size and computational complexity (Shah and Lau, 2023). This strategy offers two-fold benefits to the FedFM paradigm: primarily, it enables the participation of weak learners (clients with low-end hardware) and enhances inference efficiency by lowering memory requirements; simultaneously, it serves to alleviate communication bottlenecks by compressing the data exchanged between clients and the server. In this section, we introduce three main categories of model compression, namely federated distillation, sparsification, and quantization.
@@ 4.2. Model Compression
Table 2. Comparison of Federated Parameter-Efficient Fine-Tuning (FedPEFT) Methods.
@@
Category Representative Work Modality Model # Full Params. # Train. Params. Training Accel. Comm. Cost Selective RaFFM (Yu et al., 2023c) Txt. BERT-Large (2019) 336M 100M $6.13\times$ 29.8% FedBF (Zhang et al., 2023e) Txt. Roberta-Base (2019) 125M 0.66M $1.6\%$ Additive Adapter FedAP (Zhang et al., 2023e) Txt. Roberta-Base (2019) 125M 2M $1.6\%$ FedCLIP (Lu et al., 2023a) Vis.-Txt. ViT-B/32 (2020a) 150M 0.53M 3.5% FedDAT (Chen et al., 2024) Vis.-Txt. ALBEF (2021b) 290M 2.86M 9.9% C2A (Kim et al., 2023) Txt. DistilBERT (2020) 66M 0.06M 0.1% Fed-MNMT (Liu et al., 2023c) Txt. mBART-50 (2020) 611M 8M 1.3% AdaFL (Cai et al., 2023) Txt. BERT (2019) 110M 0.61M $1.63\times$ 0.6% Prompt PromptFL (Guo et al., 2023) Vis.-Txt. ViT-B/16 (2021) 87M 0.87M $2.38\times$ 0.9% MFPT (Zhao et al., 2024) Txt. XLM-RoBERTa (2020) 270M 1.2M 0.4% FedAPT (Su et al., 2024b) Vis.-Txt. ViT-B/32 (2020a) 88M 2.8M 3.2% FedSP (Dong et al., 2023) Txt. GPT2-XL (2019) 1.6B  [^1] M 0.5% Reparameterization-based Methods SLoRA (Babakniya et al., 2023) Txt. DistilBERT (2020) 67M 0.7M $13.47\times$ 5.8% LP-FL (Jiang et al., 2023a) Txt. BERT-Large (2019) 336M 100M 30% FedFMS (Liu et al., 2024b) Vis.-Txt. ViT-B/16 (2021) 87M 8.6M 10% pFedS2T (Du et al., 2024) Aud. Whisper (2023) 254M 10.1M 4% FFA-LoRA (Sun et al., 2024b) Txt. RoBERTa-Large (2019) 355M 0.39M 0.1%
4.2.1. Federated Distillation
As a representative mechanism for efficient model deployment, Knowledge Distillation (KD) (Hinton et al., 2015) transfers knowledge from a large teacher model to a compact student model that inherits the teacher’s capabilities by aligning their outputs (e.g., logits) or intermediate representations. In the context of FL, recent work has extended KD into a collaborative paradigm known as Federated Distillation (FD) (Jeong et al., 2018; Li et al., 2024b), where clients share compact knowledge instead of large model parameters. Depending on the direction of knowledge transfer, FD techniques can be generally categorized into two types: upstream distillation and downstream distillation, as discussed below:
Upstream Distillation.
The basic idea of this direction is to distill the knowledge of multiple local models into the shared global model. Instead of transmitting massive model parameters, clients often act as teachers (or co-learners) by uploading refined knowledge representations or lightweight adapter parameters regularized by local distillation. Several pioneering studies in this direction propose to communicate knowledge in the form of soft-label predictions on records of a public distillation (Jeong et al., 2018; Lin et al., 2020). As another representative framework, FedKD (Wu et al., 2022) employs a small model and a large model to learn and distill knowledge from each other, where only the small model is shared by different clients and learned collaboratively, hereby reduces the communication cost. Despite the high communication efficiency, these approaches fail to alleviate the computational burden, as they still require clients to optimize the full models during local training. In contrast, advanced frameworks such as FedGKT (He et al., 2020a) and FedDAT (Chen et al., 2024) explicitly train a compact module in place of the large model, reducing both communication costs and local computational burden.
Downstream Distillation.
Resource-constrained clients are often incapable of training or hosting full-sized models due to prohibitive memory and computational costs. To bridge this gap, Downstream distillation employs a pre-trained FM as a teacher to transfer its general capabilities into lightweight client-specific models. One direction involves proxy-based adaptation, such as FedPFT (Peng et al., 2024), which utilizes a proxy mechanism to facilitate the fine-tuning of FMs for clients without direct parameter access. More recent studies go beyond simple alignment to incorporate generative reasoning capabilities for language models. For instance, PPC-GPT (Fan et al., 2025d) utilizes a server-side auxiliary LLM to generate synthetic data and Chain-of-Thought (CoT) rationales based on differentially private client prompts. By distilling both the predictive labels and the reasoning processes from the FM, the server compresses the model into a task-specific SLM before deploying it to clients for local retraining. Similarly, FedCoT (Fan et al., 2025b) and other synthetic-data-driven methods explore transferring reasoning capabilities to edge devices without transmitting raw private data.
4.2.2. Sparsification
The goal of sparsification is to reduce the large number of parameters and activations in FMs, thereby mitigating communication bottlenecks and alleviating resource constraints in both training and deployment environments. In this part, we discuss sparsification techniques along three dimensions: sparsification target, phase, and granularity.
Sparsification Target.
Based on the specific object being compressed, techniques are taxonomized into three groups: (1) Communication sparsification aims to compress model updates transmitted between clients and the server, typically without altering the local training process (Jiang et al., 2023c). Representative approaches include Top-K pruning (Kuo et al., 2024; Qiu et al., 2025), which retains only updates with the largest magnitudes; thresholding (Chu et al., 2024), which filters out parameters below a specific limit; and sketching (Rothchild et al., 2020), which utilizes hashing to project updates into lower-dimensional structures. Building on these, FLASC (Kuo et al., 2024) applies sparsity to LoRA adapters, while MetaSend (Chu et al., 2024) advances thresholding by introducing a learnable module to select informative parameters adaptively. (2) Parameter sparsification targets the reduction of the model’s computational and memory footprint by pruning less critical weights, thereby enabling sparse local training. Unlike communication sparsification, this modifies the local optimization loop. For instance, FedSpaLLM (Bai et al., 2025) facilitates federated pruning, allowing clients to autonomously prune layer subsets aligned with their hardware capacities. (3) Activation sparsification enhances efficiency by selectively engaging only a subset of neurons or modules during the forward pass. This concept is mainly realized in the Mixture-of-Experts (MoE) paradigm (Jacobs et al., 1991), which utilizes a learned gating mechanism to activate a sparse set of expert subnetworks for each input.
Sparsification Phase.
While existing efforts span the entire model lifecycle, recent studies predominantly focus on training-oriented sparsification. This paradigm encompasses the aforementioned communication and parameter sparsification techniques, which operate within the training loop to accelerate convergence and reduce overheads. Some approaches also incorporate sparsification-aware learning mechanisms, such as the selective fine-tuning methods discussed in Section 4.1.1. In addition, inference-oriented sparsification targets the deployment stage, typically compressing the final global model to minimize latency and memory footprint on resource-limited devices. Research in this vein often leverages the LTH (Frankle and Carbin, 2019; Seo et al., 2021; Li et al., 2021a; Tamirisa et al., 2024) to identify accurate sparse subnetworks.
Sparsification Granularity.
Categorized by the pattern of parameter elimination, techniques fall into two groups: Unstructured sparsification involves removing individual elements (weights or gradients) without a fixed pattern. Most communication sparsification methods (e.g., Top-$k$ update compression mentioned above (Beitollahi et al., 2022; Qiu et al., 2025)) fall into this category. While effective for reducing transmission data, unstructured sparsity is often difficult to accelerate on standard hardware, failing to alleviate local computational burdens. In contrast, Structured sparsification enforces sparsity at a coarser granularity, such as removing entire channels, layers, or blocks (as seen in parameter sparsification methods like FedSpaLLM (Bai et al., 2025)). This characteristic facilitates valid sparse local training by making only specific dense substructures trainable, effectively reducing both memory usage and computational cost without requiring specialized hardware support.
4.2.3. Quantization
@@
Quantization involves reducing the precision of model parameters to decrease memory usage and computational requirements. This technique is well-established in both the FM and FL domains (Xu et al., 2024b; Reisizadeh et al., 2020). Quantization is orthogonal to other resource-efficient techniques, making it feasible to combine them for greater efficiency and flexibility (Lit et al., 2022).
Hybrid compression techniques combine multiple compression methods to achieve better performance and efficiency. For example, a hybrid approach may use both quantization and sparsification to reduce model size and improve communication efficiency simultaneously. By leveraging the strengths of different techniques, hybrid compression can lead to more effective resource utilization in federated learning settings.
4.3. Zeroth-Order Optimization